what polynomial with a degree of 4 has zeros with multiplicity 1 at x=-2 and x=2 and a zero with multiplicity 2 at x=0 and f(1)=2
Answer by ewatrrr(10682)

 
Hi,
polynomial with a degree of 4 has zeros at x=-2, x= 2 and x=0 (two)
f(x) = a(x+2)(x-2)(x-0)(x-0)
f(x) = ax^2( x^2 - 4)
f(1) = -3a = 2
a = (-2/3)
f(x) = (-2/3) x^2( x^2 - 4)